OpenStudy (anonymous):

A particular model of walkie-talkie can broadcast in a circular area. The radius of the broadcast area is 10,000 feet. Find the area of this circle to the nearest square foot. Use 3.14 for pi

OpenStudy (anonymous):

The area of a circle is equal to the radius squared times pi: \[A = \pi r^2 = 3.14 (10000 ft)^2 = 314000000 ft^2\]